# Equality of Hölder exponents for distribution functions of Gibbs measures

### v1

@@ Abstract.
We show that the pointwise Hölder spectrum of the distribution function of a Gibbs measure supported on the attractor of a self-conformal subset of the real line satisfies the multifractal formalism.
@@ 1. Introduction
(2) $\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
@@
(3) $\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Clearly, $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but the reverse inequality may fail in general. For example, if $f(t)=t^{2}$, then $\alpha_{f}(0)=\infty$ (since one can take $P(t)=t^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$. In addition, a desirable property of Hölder exponents is that they are left unchanged upon perturbation of $f$ by a smooth function $g$. Indeed, $\alpha_{f}(t)=\alpha_{f+g}(t)$, whereas $\tilde{\alpha}_{f}(t)\neq\tilde{\alpha}_{f+g}(t)$ in general.
On the other hand, for certain functions with fractal properties, the two definitions of Hölder exponent coincide. This was shown to be the case by the first author in [1, 2] for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition. In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see the next section for a precise statement.
A function $f$ is multifractal if its pointwise Hölder exponents $\alpha_{f}(t)$ take on more than one value as $t$ ranges through the domain of $f$. For such functions, we are interested in the sets
$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\}.$
The Hausdorff and packing multifractal spectra of $f$ are given, respectively, by
$\alpha\mapsto\dim_{H}E_{f}(\alpha)\qquad\mbox{and}\qquad\alpha\mapsto\dim_{P}E_{f}(\alpha),$
where $\dim_{H}$ and $\dim_{P}$ denote Hausdorff and packing dimensions, respectively. If $f=F_{\mu}$ is the distribution function of a Borel probability measure $\mu$, i.e. $F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [13, Lemma 5.1] we have
(4) $\tilde{\alpha}_{F_{\mu}}(t_{0})=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t_{0},r)\right)}{\log r}.$
For the remainder of this introduction, let $\mu$ a Gibbs measure associated with an arbitrary Hölder potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition - see the next section for precise definitions. It is well known that the sets
@@
satisfy the multifractal formalism; that is, $\dim_{H}\mathscr{E}(\alpha;\mu)=\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$, where $\beta^{*}(\alpha)$ is the Legendre transform of a suitable scaling function $\beta(q)$, defined in the next section. The second author and Sumi [13] showed recently that, in the case when all of the maps in the IFS are strictly increasing, the Hausdorff spectrum of the modified level sets
@@
In a forthcoming paper, we will show that the situation is different for the packing spectrum.
@@ 2. Notation and definitions
Let $X\subset\mathbb{R}$ be a compact interval. A $\mathcal{C}^{1+\epsilon}$ conformal iterated function system on $X$ is given by an index set $I$ with $2\leq\#I<\infty$ and $\mathcal{C}^{1+\epsilon}$ conformal contracting maps $\Phi=(\phi_{i})_{i\in I}$, $\phi_{i}:X\rightarrow X$.
More precisely, we mean that each $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$ and for each $i\in I$. For ease of presentation, we assume that each of the maps $\phi_{i},i\in I$ is strictly increasing. However, this restriction is not essential. We refer to [18] for further details and basic properties of conformal iterated function systems. Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}$ given by
@@
Let $\Lambda:=\pi(I^{\mathbb{N}})$; then $\Lambda$ satisfies the set equation $\Lambda=\bigcup_{i\in I}\phi_{i}(\Lambda)$. We refer to $\Lambda$ as the self-conformal set generated by $\Phi$. We say that $\Phi$ satisfies the open set condition if $\phi_{i}(\mathrm{Int}(X))\cap\phi_{j}(\mathrm{Int}(X))=\varnothing$ for all $i,j\in I$ with $i\neq j$. To utilize the symbolic thermodynamic formalism (see [6]) we will also need the following definitions. We denote by $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ the left-shift on $I^{\mathbb{N}}$ which becomes a compact metric space endowed with the shift metric. Let $\varphi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be the geometric potential of $\Phi$, given by
(5) $\varphi(\omega):=\log\left(\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right),\quad\omega=(\omega_{1},\omega_{2},\dots)\in I^{\mathbb{N}}.$
Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable, it follows that $\varphi$ is Hölder continuous. Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
@@
its ergodic sums. We say that a Borel probability measure $\mu_{\psi}$ on $I^{\mathbb{N}}$ is a Gibbs measure for $\psi$ if there exist two constants $C\geq 1$ and $P\in\mathbb{R}$ such that for each $n\in\mathbb{N}$ and for all $w\in I^{n}$ we have
(6) $C^{-1}\leq\frac{\mu_{\psi}[w_{1},\dots,w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
@@
We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
(7) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{x,y\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<+\infty.$
@@
It is well known that the multifractal spectrum $\mathcal{F}(\alpha)$ is complete [28], that is, $\mathcal{F}(\alpha)\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$. Moreover, the following dichotomy is well known [25]: either we have $\alpha_{-}<\alpha_{+}$, or $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$. Note that in the latter case we have $\alpha_{-}=\alpha_{+}=\dim_{H}\Lambda$.
Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)$ such that
@@ 3. The main theorem
Let $\Phi=(\phi_{i}:X\rightarrow X)_{i\in I}$ be a conformal iterated function system satisfying the open set condition. Let $\varphi$ denote the geometric potential of $\Phi$ defined in (5) and let $\psi$ be any Hölder potential with $\mathcal{P}(\psi)=0$ such that $\psi$ is not cohomologous to $\varphi$. Let $\mu_{\psi}$ be a Gibbs measure for $\psi$. Then, for the distribution function $F$ of $\mu_{\psi}\circ\pi^{-1}$, we have that $\alpha_{F}=\tilde{\alpha}_{F}$.
@@
The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out some degenerate cases. As a simple example, consider the IFS $\Phi=\{x/2,(x+1)/2\}$, whose attractor is the interval $[0,1]$. Clearly, $\varphi\equiv-\log 2$. Choosing $\psi\equiv-\log 2$ as well, we see that $\mu:=\mu_{\psi}\circ\pi^{-1}$ is Lebesgue measure on $[0,1]$, so $F$ is infinitely differentiable on $(0,1)$ and hence $\alpha_{F}(t)\equiv\infty$ on $(0,1)$. On the other hand, it is easy to see that $\tilde{\alpha}_{F}(t)\equiv 1$ on $(0,1)$.
More generally, we claim that for a Hölder continuous $\psi$ with $\mathcal{P}(\psi)=0$ we have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some/any Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity. To prove this, first assume that $\psi$ is cohomologous to $\varphi$. Hence, $P(\varphi)=P(\psi)=0$ and thus, $\dim_{H}(\Lambda)=1$. Hence, by the well-known construction of conformal measures (see e.g. [19, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure. Hence, by Lebesgue’s density theorem, we conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and thus, $\Lambda=X$ is an interval. Indeed, if $X\setminus\bigcup_{i\in I}\Phi_{i}(X)$ was non-empty, it would contain an interval of positive length. By standard arguments, this would imply that $\Lambda$ contains no Lebesgue density points, which gives the desired contradiction. Consequently, the normalised Lebesgues measure on $\Lambda$ is a Gibbs measure for $\varphi$. Since $\psi$ is cohomologous to $\varphi$, $\mu_{\psi}$ is also a Gibbs measure for $\varphi$. It follows that $\mu_{\psi}$ is equivalent to the Lebesgue measure with a density bounded away from zero and infinity.
Conversely, assume that $\Lambda$ is an interval and $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to Lebesgue with density bounded away from zero and infinity. Hence, Lebesgue measure is also Gibbs measure for $\psi$. Since Lebesgue measure is also a Gibbs measure for $\varphi$ and $P(\varphi)=P(\psi)=0$, it follows from [19, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$.
The key to proving Theorem 3.1 is to prove Proposition 3.4 below. The proof of that proposition uses the following lemma.

### v2

@@ Abstract.
Pointwise Hölder exponents describe the degree of regularity of a function near a point. For a function $f:\mathbb{R}\to\mathbb{R}$, a number $\alpha>0$ and a point $t_{0}\in\mathbb{R}$, write $f\in C^{\alpha}(t_{0})$ if there is a constant $C$ and a polynomial $P$ of degree less than $\alpha$ such that
$|f(t)-P(t-t_{0})|\leq C|t-t_{0}|^{\alpha}\qquad\mbox{for all $t\in\mathbb{R}$}.$
The pointwise Hölder exponent of $f$ at $t_{0}$ is the number
$\alpha_{f}(t_{0}):=\sup\{\alpha>0:f\in C^{\alpha}(t_{0})\}.$
A simpler quantity, also frequently called pointwise Hölder exponent in the mathematical literature, is the number
$\tilde{\alpha}_{f}(t_{0}):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t_{0})\},$
where $f\in\tilde{C}^{\alpha}(t_{0})$ means that there is a constant $C>0$ such that $|f(t)-f(t_{0})|\leq C|t-t_{0}|^{\alpha}$ for all $t\in\mathbb{R}$. Clearly $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but strict inequality is possible and in fact common. In this paper we consider the case when $f=F_{\mu}$ is the distribution function of a Gibbs measure $\mu$ associated with an arbitrary Hölder continuous potential $\psi$ on a self-conformal set, and show that, under a very mild condition on $\psi$, $\alpha_{f}(t)=\tilde{\alpha}_{f}(t)$ for all $t$. As a consequence, we deduce that the pointwise Hölder spectrum of $f$ satisfies the multifractal formalism. As an application, we derive the pointwise Hölder spectrum of conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps of an interval.
Key words and phrases:
Hölder exponent, Gibbs measure, Conformal iterated function system, Distribution function, Multifractal spectrum, Bounded distortion property
@@ 1. Introduction
$\alpha_{f}(t):=\sup\{\alpha>0:f\in C^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Of great interest are the level sets
$E_{f}(\alpha):=\{t\in\mathbb{R}:\alpha_{f}(t)=\alpha\},\qquad\alpha\geq 0.$
The multifractal spectrum, or pointwise Hölder spectrum, of $f$ is the function
$\alpha\mapsto\dim_{H}E_{f}(\alpha),$
where $\dim_{H}$ denotes Hausdorff dimension.
Hölder spectra are an important analytical tool in the study of certain physical processes that exhibit a wide range of local regularity behavior. They were introduced by Mandelbrot [15] and Frisch and Parisi [8] to study intermittent turbulence flows and intensity of seismic waves. Jaffard [10, 11] determined Hölder spectra for a class of self-similar functions satisfying a certain smoothness condition from a mathematical point of view using wavelet theory. Subsequent authors [4, 5, 12, 23] computed Hölder spectra of specific self-affine functions; their results were generalized by the first author in [1, 2]. For pointwise Hölder spectra of Weierstrass-type functions, we refer to [18].
@@
$\tilde{\alpha}_{f}(t):=\sup\{\alpha>0:f\in\tilde{C}^{\alpha}(t)\},\qquad t\in\mathbb{R}.$
Clearly, $\alpha_{f}(t)\geq\tilde{\alpha}_{f}(t)$, but the reverse inequality may fail in general. For example, if $f(t)=t^{2}$, then $\alpha_{f}(t_{0})=\infty$ for all $t_{0}\in\mathbb{R}$ (since one can take $P(t)=(t+t_{0})^{2}$ in (1) for all $\alpha>2$), but $\tilde{\alpha}_{f}(0)=2$ whereas $\tilde{\alpha}_{f}(t_{0})=1$ for all $t_{0}\neq 0$. In addition, a desirable property of Hölder exponents is that they are left unchanged upon perturbation of $f$ by a smooth function $g$. Indeed, $\alpha_{f}(t)=\alpha_{f+g}(t)$, whereas $\tilde{\alpha}_{f}(t)\neq\tilde{\alpha}_{f+g}(t)$ in general.
Unfortunately, the two definitions of pointwise Hölder exponent have frequently been conflated in the literature, often without proper justification. For this reason we feel that it is critically important to study classes of functions, ideally as general as possible, for which the two notions of Hölder exponent coincide.
In general, if $f\in C^{\alpha}(t_{0})$ for $\alpha>1$ and $f^{\prime}(t_{0})\neq 0$, then $\tilde{\alpha}_{f}(t_{0})=1<\alpha_{f}(t_{0})$. Turning this around, it follows that if $\alpha_{f}(t_{0})=\tilde{\alpha}_{f}(t_{0})$, then either $\alpha_{f}(t_{0})\leq 1$ or $f^{\prime}(t_{0})=0$. This shows that the two definitions of Hölder exponent disagree for most functions known from calculus. On the other hand, many sufficiently irregular functions have the property that the only possible finite value of their derivative is zero. For example, it was shown by the first author in [1, 2] that for certain classes of self-affine functions, including distribution functions of self-similar measures on $\mathbb{R}$ satisfying the open set condition, the two definitions of pointwise Hölder exponent coincide. In this note, we extend this result to distribution functions of arbitrary Gibbs measures on finitely generated self-conformal sets; see Theorem 3.1 for a precise statement. Although it may seem plausible that the result can be extended from self-similar measures to self-conformal ones, the proof nonetheless requires an entirely new approach. This is because the method of proof in [2], using divided differences, depends very explicitly on the strict self-similarity of the measure and can therefore not be extended to the self-conformal setting.
If $f=F_{\mu}$ is the distribution function of a Borel probability measure $\mu$, i.e. $F_{\mu}(x):=\mu\left((-\infty,x]\right)$ for $x\in\mathbb{R}$, then by [9, Lemma 5.1] we have for each $t\in\mathbb{R}$,
(2) $\tilde{\alpha}_{F_{\mu}}(t)=\liminf_{r\rightarrow 0}\frac{\log\mu\left(B(t,r)\right)}{\log r}.$
For the remainder of this introduction, let $\mu$ be a Gibbs measure associated with an arbitrary Hölder continuous potential on the attractor of a conformal iterated function system (IFS) satisfying the open set condition – see the next section for precise definitions. It is well known that the sets
@@
satisfy the multifractal formalism; that is, $\dim_{H}\mathscr{E}(\alpha;\mu)=\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$, where $\beta^{*}(\alpha)$ is the Legendre transform of a suitable scaling function $\beta(q)$, defined in the next section. This result was first proved in [19] for conformal iterated functions, and in [21] for conformal repellers (see also [20] for a detailed account and related results). The second author and Sumi [9] showed recently that, in the case when all of the maps in the IFS are strictly increasing, the Hausdorff spectrum of the modified level sets
@@
In a forthcoming paper, we will show that the situation is different for the packing spectrum, that is, the function $\alpha\mapsto\dim_{P}E_{f}(\alpha)$, where $\dim_{P}$ denotes the packing dimension.
As an application of our main theorem, we determine the pointwise Hölder spectrum of conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps in Corollary 5.1. This result complements [9, Corollary 2.2], where the dimension spectrum of the conjugacy map is based on the simpler definition of the Hölder exponent $\tilde{\alpha}$. We note that it is straightforward to extend Corollary 5.1 to the framework of expanding circle endomorphisms. For related results on the non-differentiability of conjugacy maps in both frameworks we refer to [13].
Remark 1.1.
A closely related concept is that of $\alpha$-Hölder differentiability. For $\alpha>0$, say a function $f$ is $\alpha$-Hölder differentiable at a point $t_{0}$ if the limit
$\lim_{t\to t_{0}}\frac{|f(t)-f(t_{0})|}{|t-t_{0}|^{\alpha}}$
exists. The existence of the above limit (0, positive and finite, or $\infty$) gives information about the Hölder exponent $\tilde{\alpha}_{f}(t_{0})$ (but not, in general, about $\alpha_{f}(t_{0})$), and vice versa. Hölder differentiability for distribution functions of Gibbs measures was considered in [7, 14], with a later improvement in [24]. Although some of the methods used in these papers (including the use of pressure functions and the bounded distortion principle) are similar to those used here, the principal questions are fundamentally different.
This paper is organized as follows: In Section 2, we introduce the relevant notation and definitions. In Section 3, we state and prove our main result, namely the equality of $\alpha_{F}(t)$ and $\tilde{\alpha}_{F}(t)$ when $F$ is the distribution function of a Gibbs measure associated with a Hölder continuous potential $\psi$ on a self-conformal attractor. The only condition we impose on $\psi$ is that it is not cohomologous to the geometric potential $\varphi$ of the IFS; this is to avoid cases where $F$ is absolutely continuous and hence too “regular”. The technical heart of the proof is Proposition 3.7, which states that for each positive integer $k$ the limit
$\lim_{y\to x}\frac{F(y)-F(x)}{(y-x)^{k}},$
if it exists, cannot be nonzero and finite.
As a consequence of the main result, we derive in Section 4 that the (Hausdorff) multifractal spectrum of such a distribution function is given by the multifractal formalism. In this section we show also that, at least in the self-similar case, the packing multifractal spectrum does not satisfy the multifractal formalism. Finally, in Section 5, we give an application of our main theorem to conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps.
@@ 2. Notation and definitions
Let $X\subset\mathbb{R}$ be a compact interval of positive length. A $\mathcal{C}^{1+\epsilon}$ conformal iterated function system on $X$ is given by an index set $I$ with $2\leq\#I<\infty$ and $\mathcal{C}^{1+\epsilon}$ conformal contracting maps $\Phi=(\phi_{i})_{i\in I}$, $\phi_{i}:X\rightarrow X$. More precisely, we mean that, for each $i\in I$, the map $\phi_{i}$ extends to a $\mathcal{C}^{1+\epsilon}$ differentiable map on a neighborhood of $X$ which satisfies $0<\left|\phi_{i}^{\prime}(x)\right|<1$ for every $x\in X$. We refer to [16] for further details and basic properties of conformal iterated function systems. Let $\pi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ denote the coding map of $\Phi$ which is for $\omega=(\omega_{1}\omega_{2}\dots)\in I^{\mathbb{N}}$ given by
@@
Let $\Lambda:=\pi(I^{\mathbb{N}})$; then $\Lambda$ satisfies the set equation $\Lambda=\bigcup_{i\in I}\phi_{i}(\Lambda)$. We refer to $\Lambda$ as the self-conformal set generated by $\Phi$. We say that $\Phi$ satisfies the open set condition if $\phi_{i}(\mathrm{Int}(X))\cap\phi_{j}(\mathrm{Int}(X))=\varnothing$ for all $i,j\in I$ with $i\neq j$. To utilize the symbolic thermodynamic formalism (see [6]) we will also need the following definitions. We denote by $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ the left-shift on $I^{\mathbb{N}}$ which becomes a compact metric space endowed with the shift metric. A basis for the topology is given by the cylinder sets
$w_{1}\dots w_{n}]:=\{\omega\in I^{\mathbb{N}}:\omega_{1}=w_{1},\dots,\omega_{n}=w_{n}\},\qquad w_{1}\dots w_{n}\in I^{n},\quad n\in\mathbb{N}.$
Let $\varphi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be the geometric potential of $\Phi$, given by
(3) $\varphi(\omega):=\log\left|\phi_{\omega_{1}}^{\prime}(\pi(\sigma(\omega)))\right|,\quad\omega=\omega_{1}\omega_{2}\dots\in I^{\mathbb{N}}.$
Since each $\phi_{i}$ is $\mathcal{C}^{1+\epsilon}$ differentiable and the maps $\phi_{i}$ are uniformly contracting, it follows that $\varphi$ is Hölder continuous. Let $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ be Hölder continuous and denote by
@@
its ergodic sums. We will frequently make use of the fact that any Hölder continuous $\psi:I^{\mathbb{N}}\rightarrow\mathbb{R}$ satisfies the bounded distortion property
(4) $\sup_{n\in\mathbb{N}}\sup_{w\in I^{n}}\sup_{\rho,\tau\in I^{\mathbb{N}}}\left|S_{n}\psi(w\rho)-S_{n}\psi(w\tau)\right|<\infty.$
We say that a Borel probability measure $\mu_{\psi}$ on $I^{\mathbb{N}}$ is a Gibbs measure for $\psi$ if there exist two constants $C\geq 1$ and $P\in\mathbb{R}$ such that for each $n\in\mathbb{N}$ and for all $w\in I^{n}$ we have
(5) $C^{-1}\leq\frac{\mu_{\psi}[w_{1}\dots w_{n}]}{e^{S_{n}\psi(\overline{w})-nP}}\leq C.$
@@
Two Hölder potentials $\psi$ and $\xi$ on $I^{\mathbb{N}}$ are said to be cohomologous if there is a bounded potential $\eta:I^{\mathbb{N}}\to\mathbb{R}$ such that $\psi-\xi=\eta-\eta\circ\sigma$. It is easy to see that, if $\psi$ and $\xi$ are cohomologous, then $\mathcal{P}(\psi)=\mathcal{P}(\xi)$.
@@
It is well known that the multifractal spectrum is complete [22], that is, $\mathscr{E}(\alpha;\mu_{\psi})\neq\varnothing$ if and only if $\alpha\in\left[\alpha_{-},\alpha_{+}\right]$. Moreover, the following dichotomy is well known [21]: either we have $\alpha_{-}<\alpha_{+}$, or $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$. Note that in the latter case we have $\alpha_{-}=\alpha_{+}=\dim_{H}\Lambda$.
Since $\varphi<0$, there is for each $q\in\mathbb{R}$ a unique $\beta(q)\in\mathbb{R}$ such that
@@ 3. The main theorem
Let $\Phi=(\phi_{i}:X\rightarrow X)_{i\in I}$ be a conformal iterated function system satisfying the open set condition. Let $\pi:I^{\mathbb{N}}\to\mathbb{R}$ denote the coding map of $\Phi$. Let $\varphi$ denote the geometric potential of $\Phi$ defined in (3) and let $\psi$ be any Hölder potential with $\mathcal{P}(\psi)=0$ such that $\psi$ is not cohomologous to $\varphi$. Let $\mu_{\psi}$ be a Gibbs measure for $\psi$. Then, for the distribution function $F$ of $\mu_{\psi}\circ\pi^{-1}$, we have that $\alpha_{F}=\tilde{\alpha}_{F}$.
@@
The condition that $\psi$ is not cohomologous to $\varphi$ is needed to rule out cases where $F$ is too “smooth”. As a simple example, consider the IFS $\Phi=\{x/2,(x+1)/2\}$, whose attractor is the interval $[0,1]$. Clearly, $\varphi\equiv-\log 2$. Choosing $\psi\equiv-\log 2$ as well, we see that $\mu:=\mu_{\psi}\circ\pi^{-1}$ is Lebesgue measure on $[0,1]$, so $F$ is infinitely differentiable on $(0,1)$ and hence $\alpha_{F}(t)\equiv\infty$ on $(0,1)$. On the other hand, it is easy to see that $\tilde{\alpha}_{F}(t)\equiv 1$ on $(0,1)$.
More generally, we have the following characterization:
Proposition 3.3.
We have that $\psi$ is cohomologous to $\varphi$ if and only if $\Lambda$ is an interval and some (hence any) Gibbs measure $\mu_{\psi}$ for $\psi$ satisfies that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to the Lebesgue measure with a Radon-Nykodym density bounded away from zero and infinity.
Proof.
First assume that $\psi$ is cohomologous to $\varphi$. Hence, $P(\varphi)=P(\psi)=0$ and thus, by Bowen’s formula (see e.g. [17, Theorem 4.2.11]), $\Lambda$ has positive Lebesgue measure. We conclude that $X=\bigcup_{i\in I}\Phi_{i}(X)$ and therefore, $\Lambda=X$ is an interval. Indeed, if $X\setminus\bigcup_{i\in I}\Phi_{i}(X)$ was non-empty, it would contain an interval of positive length. By standard arguments, this would imply that $\Lambda$ contains no Lebesgue density points, which gives the desired contradiction. Let $\lambda$ denote the normalised Lebesgue measure on $X$. Since $\Lambda=X$ and $P(\varphi)=0$, the unique Borel probability measure $\mu$ on $I^{\mathbb{N}}$ satisfying $\mu\circ\pi^{-1}=\lambda$ is a Gibbs measure for $\varphi$. Since $\psi$ is cohomologous to $\varphi$, it follows from [17, Theorem 2.2.7] that any Gibbs measure $\mu_{\psi}$ for $\psi$ is equivalent to $\mu$ with a density bounded away from zero and infinity. Hence, $\mu_{\psi}\circ\pi^{-1}$ is equivalent to $\lambda$ with a density bounded away from zero and infinity.
Conversely, assume that $\Lambda$ is an interval. Then, $\Lambda=X$ and we have $P(\varphi)=0$ by Bowen’s formula. Further, the unique Borel probability measure $\mu$ on $I^{\mathbb{N}}$ defined by $\mu\circ\pi^{-1}=\lambda,$ is a Gibbs measure for $\varphi$. Assume that $\mu_{\psi}$ is a Gibbs measure for $\psi$ satisfying that $\mu_{\psi}\circ\pi^{-1}$ is absolutely continuous with respect to $\lambda$ with density bounded away from zero and infinity. Then, $\mu$ is also a Gibbs measure for $\psi$. Since $P(\varphi)=P(\psi$), it follows from [17, Theorem 2.2.7] that $\varphi$ is cohomologous to $\psi$. ∎
The remainder of this section is devoted to the proof of Theorem 3.1. We will assume without loss of generality that the images of $X$ under the maps $\phi_{1},\dots,\phi_{m}$ are arranged from left to right; that is, $\inf\phi_{i+1}(X)\geq\sup\phi_{i}(X)$ for $i=1,\dots,m-1$, where $m:=|I|$. The first lemma is needed to deal with the codings of points that lie close to the common endpoint of adjacent cylinder intervals. The precise conclusion depends on the orientations of the first and last maps, $\phi_{1}$ and $\phi_{m}$, of the IFS. In this section, we write $\mathbf{i}|_{n}:=i_{1}\dots i_{n}$ for $\mathbf{i}=i_{1}\dots i_{p}\in I^{p}$ with $p\geq n$; and similarly, $\omega|_{n}:=\omega_{1}\dots\omega_{n}$ for $\omega=\omega_{1}\omega_{2}\dots\in I^{\mathbb{N}}$.